Q Search this course 1 Homework signment: Module 1 Homework aveSubmit Assignmant for Gra eestions Prblem 4.1 Check My Work (3 remaining Cick here to read the eBook: Profitablity Ratios Problem Walk-Through is 0 RATIO CALCULATIONS thomson Trucking has S22 bilion in assets, and its tax rate is 40% lts basc earning power (BEP) abo is 20%, and n·return onessets (ROA) a 3% what is its times-in etwas me ate, Round your answer to two decimal places. 20.
Answer #1
Return on assets: 3%
Total assets : 22 billion or 22000 millions
Return on assets = Net income after tax/ Total assets *100
3   = Net income after tax / 22000 *100
Net Income after tax = 22000*3/100 = 660 million
Earnings after interest before tax:
Net income after tax: 660
Add: Tax (660/60*40) 440
Earnings after interest before tax: 1100
basic earning power =20%
Basic Earnings power = EBIT/ Total assets *100
20 = EBIT /22000 *100
EBIT = 4400 million
Interest expenses: EBIT -Earnings after interest before tax
4400- 1100 = 3300 millions
Times interest earned = EBIT / Interest expense
4400/ 3300 = 1.33 times